OpenStudy (anonymous):

If x is real and p = (3(x^2 + 1) / (2x - 1) , prove that p^2 - 3(p + 3) >= 0

OpenStudy (jack1):

if: p = (3(x^2 + 1) / (2x - 1) therefore p^2 - 3(p + 3) = (9 ( -x^2 + x + 1)^2)/(1 - 2x)^2

OpenStudy (jack1):

= (9x^4 - 18x^3 - 9x^2 + 18x + 9) --------------------------------- (4x^2 - 4x + 1)

OpenStudy (anonymous):

\[p^2 - 3(p +3) >= 0\]

OpenStudy (anonymous):

\[p = (3(x^2 + 1) / (2x -1))\]

OpenStudy (jack1):

sure, so when does \[\frac{ (9x^4 - 18x^3 - 9x^2 + 18x + 9) }{ (4x^2 - 4x + 1) } = 0or +ve\]

OpenStudy (anonymous):

how did you do that?

OpenStudy (anonymous):

because p=(3(x2+1)/(2x−1))^2 gives \[((9x^4 +18x^2 +9) / (4x^2 - 4x + 1))\]

OpenStudy (jack1):

u havent added in -3(p + 3)

OpenStudy (anonymous):

and how you do that

OpenStudy (jack1):

\[\frac{ (9x+18x ) } { (2 x-1) }\]

OpenStudy (jack1):

\[\frac{ 9x^2 +18x }{ 2x -1 }\times \frac{ 2x -1 }{2x -1 }\]

OpenStudy (anonymous):

solutin pls im still confused?

OpenStudy (jack1):

\[= \frac{ 18 x^3+27 x^2-18 x }{4x^2 - 4x +1 }\]

OpenStudy (jack1):

as the 2 equations now have the same denominator, u can add the numerators together

OpenStudy (jack1):

\[(9x^4 +18x^2 + 9) - (18x^3+27x^2−18x)\]

OpenStudy (jack1):

\[ = (9x^4 -18x^3 +18x^2 - 27x^2 + 18x + 9) \]